OpenStudy (anonymous):

HOw many solutions does this equation have? (9/x-5)+(4/x^2-3x-10)=(6/x+2)

OpenStudy (mathteacher1729):

Put the denominator of each fraction it is own set of parenthesis, please. In other words, please write 1 / (x+1) instead of 1 / x + 1 :)

OpenStudy (anonymous):

HOw many solutions does this equation have? 9/(x-5)+4/(x^2-3x-10)=6/(x+2)

OpenStudy (mathteacher1729):

Ok , that is much better. :) Thank you. First question -- can you factor (x^2 - 3x -10) ?

OpenStudy (anonymous):

yeah (x-3)(x+5) right?

OpenStudy (anonymous):

wait nevermind

OpenStudy (anonymous):

(x-5)(x+2)

OpenStudy (anonymous):

Now what do I do

OpenStudy (mathteacher1729):

Now we find common denominators.

OpenStudy (mathteacher1729):

9/(x-5)+4/(x^2-3x-10)=6/(x+2) factor the quadratic to find 9/(x-5)+4/(x-5)(x+2)=6/(x+2) now make sure everything has a (x+2)(x-5) in the denominator.

OpenStudy (anonymous):

Okay got that.

OpenStudy (anonymous):

so do I times both sides by (x-5)(x+2)

OpenStudy (anonymous):

then I get 9(x+2)+4=6(x-5) right?

OpenStudy (anonymous):

then I get 9x=6x-52 then how do i solve that

OpenStudy (anonymous):

?

OpenStudy (mathteacher1729):

>then I get 9(x+2)+4=6(x-5) right? This is correct. Now distribute the 9 and the 6.

OpenStudy (anonymous):

yes so I get, 9x=6x-52